OpenStudy (anonymous):

How to prove that \(\sum_{k=0}^{\frac{n}{2}} C(n,2k) 3^{n-2k} = 2^{n-1}(2^n+1)\)?

OpenStudy (anonymous):

@KingGeorge Would you mind giving a hand here?

OpenStudy (kinggeorge):

Presumably \(n\) is even, so let's first rewrite the sum as\[\sum_{k=0}^n\binom{2n}{2k}3^{2n-2k}\]and see if that helps.

OpenStudy (kinggeorge):

I'm just going to jot down whatever I can think of, and maybe I'll get somewhere.\[(1+3)^{2n}=\sum_{k=0}^{2n}\binom{2n}{k}3^{2n}=4^{2n}\]

OpenStudy (kinggeorge):

\[2^{2n-1}(2^{2n}+1)=2^{4n-1}+2^{2n-1}=\frac{4^{2n}}{2}+\frac{4^n}{2}\]

OpenStudy (kinggeorge):

We also have the identity\[(-3+1)^{2n}=\sum_{k=0}^{2n}\binom{2n}{k}(-3)^{2n-k}=(-2)^{2n}=2^{2n}\]

OpenStudy (kinggeorge):

And \[(3^2+1)^n=\sum_{k=0}^n\binom{n}{k}3^{2n-2k}=10^n=2^n5^n\]

OpenStudy (kinggeorge):

Ah. I think I finally got something. Consider this\[\begin{aligned} (-3+1)^{2n}+(3+1)^{2n}&=\sum_{k=0}^{2n}\binom{2n}{k}(-3)^{2n-k}+\sum_{k=0}^{2n}\binom{2n}{k}(3)^{2n-k} \\ &=\sum_{k=0}^{2n}\binom{2n}{k}\left(3^{2n-k}+(-3)^{2n-k}\right)\\ &=\sum_{k=0}^{2n}\binom{2n}{2k}2\cdot3^{2n-2k}\\ &=(-2)^{2n}+4^{2n}=2^{2n}+4^{2n}. \end{aligned}\] In the last sum I wrote, we know that \(\binom{2n}{2k}=0\) for any \(k>n\), and we can pull the 2 out to get\[2^{2n}+2^{4n}=2\sum_{k=0}^n\binom{2n}{2k}3^{2n-2k}\]and so when we divide by 2, it follows that\[\sum_{k=0}^n\binom{2n}{2k}3^{2n-2k}=2^{2n-1}+2^{4n-1}=2^{2n-1}(2^n+1).\]